OpenStudy (anonymous):

What property is illustrated by the statement? 7 + (4 + 4) = (7 + 4) + 4 A. Inverse Property of Addition B. Associative Property of Addition C. Commutative Property of Multiplication D. Commutative Property of Addition

OpenStudy (anonymous):

D

OpenStudy (anonymous):

(a+b)+c=a+(b+c) ,Associative a+b=b+a,Commutative a × (b + c) = a × b + a × c, Distributive. The answer is D :)